# Rethinking Data Augmentation for Robust LiDAR Semantic Segmentation in Adverse Weather

### v1

@@ Abstract
Existing LiDAR semantic segmentation methods often struggle with performance declines in adverse weather conditions. Previous research has addressed this issue by simulating adverse weather or employing universal data augmentation during training. However, these methods lack a detailed analysis and understanding of how adverse weather negatively affects LiDAR semantic segmentation performance. Motivated by this issue, we identified key factors of adverse weather and conducted a toy experiment to pinpoint the main causes of performance degradation: (1) Geometric perturbation due to refraction caused by fog or droplets in the air and (2) Point drop due to energy absorption and occlusions. Based on these findings, we propose new strategic data augmentation techniques. First, we introduced a Selective Jittering (SJ) that jitters points in the random range of depth (or angle) to mimic geometric perturbation. Additionally, we developed a Learnable Point Drop (LPD) to learn vulnerable erase patterns with Deep Q-Learning Network to approximate the point drop phenomenon from adverse weather conditions. Without precise weather simulation, these techniques strengthen the LiDAR semantic segmentation model by exposing it to vulnerable conditions identified by our data-centric analysis. Experimental results confirmed the suitability of the proposed data augmentation methods for enhancing robustness against adverse weather conditions. Our method attains a remarkable 39.5 mIoU on the SemanticKITTI-to-SemanticSTF benchmark, surpassing the previous state-of-the-art by over 5.4%p, tripling the improvement over the baseline compared to previous methods achieved.
@@ 1 Introduction
LiDAR semantic segmentation is a fundamental task of 3D scene understanding. Robust and reliable segmentation is crucial in safety-critical applications like autonomous driving. However, existing LiDAR segmentation models [5, 1, 15, 29, 22, 38] commonly lack robustness, reporting significant performance degradation under adverse weathers such as snow, fog, rain or wet surfaces.
@@
To address this issue, recent studies [12, 33, 34] have introduced corruption benchmarks including adverse weather and proposed effective techniques for robust LiDAR segmentation against corruption. These robust methods are primarily divided into two kinds: a task-agnostic approach and a simulation-based approach. The task-agnostic approach employs general machine-learning approaches for robustness without explicitly addressing the characteristics of LiDAR corruption caused by adverse weather. The simulation-based approach [8, 3, 24, 7, 10, 35, 27] artificially synthesizes weather-specific data through physical equations for training. However, these efforts focus on detection tasks, and each method considers only a single type of weather at a time.
@@
While the simulation-based approach can better exploit the intrinsic properties of LiDAR scans under target adverse weather, simulating every weather type in all possible severities is often impractical and inaccurate. Instead, we adopt a data-centric perspective to analyze corrupted LiDAR data. For example, the distortion caused by “rain”, “snow” and “fog” often create similar patterns, as simulated in [10]. Also, all those adverse weather shows point drop through attenuation or occlusion by droplet [10, 7, 6]. While the simulation-based approach requires explicit modeling of the LiDAR for each weather condition, our data-centric approach potentially addresses complex distortions with a few unified patterns. Through existing research and our comprehensive analysis, we find that various adverse weather conditions create similar distortion patterns in LiDAR data. From this insight, we categorize these distortions into two types: (1) geometric perturbation and (2) point drop. By conducting a toy experiment, we demonstrate that these two distortion types are highly correlated with segmentation performance degradation.
@@
Rooted by the toy experiment, we introduce two novel and strategic data augmentation methods tailored to the LiDAR distortion caused by adverse weather. By incorporating these augmentations during training, we aim to enhance the model’s robustness for each distortion type. For geometric perturbation, we introduce Selective Jittering (SJ), which applies Additive Gaussian Noise (AGN) to alter the XYZ-coordinate and intensity within a selective local region. For handling a point drop pattern, we developed a Learnable Point Drop (LPD) that employs Deep Q-Network (DQN) to strategically remove points. Our augmentation strategies are informed by existing studies [10, 17, 28], which shows geometric perturbation typically involves small, random alterations to the original points. This insight led to our choice of jittering as a suitable technique for addressing the geometric perturbation. Also, augmentation strategy for point drop or occlusion is motivated from existing studies [8, 10, 7] which state that intensity and depth of LiDAR beams play a significant role. Therefore, our idea developed that using a DQN to detect point drops that impair LiDAR semantic segmentation models might enable the DQN to effectively follow and adapt to such point drop patterns.
@@
Finally, we apply the proposed augmentations when training the LiDAR segmentation model with a clean source dataset. Experimental results demonstrate that our method achieves new state-of-the-art performance on SemanticSTF, a real adverse weather dataset. Remarkably, we achieved an impressive +8.1 mIoU gain over the baseline model. This is a notable improvement, nearly tripling the +2.5 mIoU improvement from prior work [33]. Furthermore, our method has compatibility with other training techniques (e.g., improving the model architecture and training recipes) for LiDAR segmentation models, thereby offering the potential for synergistic benefits. This compatibility, along with our significant mIoU gains, highlights the efficacy of our data augmentations.
@@ 2.1 LiDAR Semantic Segmentation
Voxel-based methods [5, 38, 15] perform efficient computations by dividing 3D space into a voxel grid and aggregating points within the same voxel. MinkUnet [5] voxelizes LiDAR points with a cubic grid and applies sparse convolution. Cylinder3D [38] proposes cylindrical partitions, reflecting LiDAR’s characteristic that the density of points depends on the distance. SphereFormer [15] utilizes a radial window and transformer architecture to aggregate long-range information and improve performance. Voxel-based methods achieve a balance between reasonable inference time and commendable segmentation performance.
@@ 2.3 LiDAR Under Adverse Weather Conditions
Robustness under harsh conditions is crucial in safety-critical applications. Adverse weather substantially degrade performance in real-world outdoor autonomous driving. Thus, there are several attempts to develop weather-robust models in fields such as 2D segmentation [16, 18], 3D detection [9, 7, 12], and 3D segmentation [33, 12]. Simulation-based approaches [9, 7] artificially synthesize data for single weather conditions through physical modeling and utilize it for training. We differ from these in that we do not model specific weather conditions explicitly. Recently proposed task-agnostic approaches [33, 12] consider multiple weather conditions at once. However, they use general machine-learning techniques (such as teacher-student framework and feature prototype) to achieve robustness rather than specifically tackle LiDAR corruption caused by adverse weather. We differ from these in that we propose augmentation methods specifically tailored for adverse weather conditions based on the analysis of performance degradation in LiDAR data.
@@ Methods
Based on previous research findings mentioned in Section 3, the types of distortions in LiDAR point clouds caused by adverse weather conditions converge into a set of common distortions. Therefore, from a data-centric perspective, the issues that we need to consider can be summarized as follows: (D1) Point drop, (D2) Occlusions, (D3) Geometric perturbation, and (D4) Intensity distortion. Here, our aim is to identify which distortion types negatively impact performance. To achieve this, we have generated four distortion types of toy synthetic data through the SemanticKITTI validation set.
@@ 4 Methods
Figure 2: Illustration of proposed selective jittering. (a) Original points from clean data. (b) Depth-selective jittering (DSJ) adds Gaussian noise to a random range of depths. (c) Angle-selective jittering (ASJ) adds Gaussian noise to a random range of angles. (d) Range Jittering adds Gaussian noise in the range direction and is used alongside points treated with DSJ or ASJ.
@@ 4.1 Selective Jittering
Depth-Selective Jittering (DSJ). DSJ employs a single scan to add Gaussian noise to the XYZ-coordinates and intensity value of points below a randomly selected depth range.
@@ 4.2 Learnable Point Drop
In Comparison with random drop, random drop uniformly reduces points across all depths. Consequently, this approach is not suitable for simulating point drops caused by fog, as it lacks the depth-specific characteristics of fog-related distortions. Furthermore, randomly generated point drop scans may not achieve the severity of adverse weather conditions necessary to drop critical points that could deceive the model. Detailed comparison will be conducted in the ablation study section.
@@ 5.1 Experimental setup
Dataset and Evaluation Metrics. To assess robustness, we trained our model on 19 classes from SemanticKITTI and then validated its performance on validation set of SemanticSTF as same as [33]. The experiment process involves training on clean data followed by evaluating robustness in adverse conditions, akin to existing research [12, 33, 34]. SemanticSTF is a dataset gathered under real-world adverse conditions such as rain, fog, and snow. Following the approach in [33], invalid data in SemanticSTF are mapped to an “ignore label”. We utilized MinkowskiNet [5] for validation purposes due to its well-known robustness [12, 34]. Our evaluation metrics include the Intersection over Union (IoU) for each class and the mean IoU (mIoU) across all classes. Experiments on synthetic data like SynLiDAR are detailed in the supplementary material due to its low significance because our experiment targets high performance in adverse weather using real data.
@@ 5.2 Main results
Table 2: LiDAR segmentation results (mIoU) on the SemanticSTF validation set of models trained with SemanticKITTI. D-fog and L-fog denote dense fog and light fog weather condition in all experiments. is the reproduced result with same segmentation backbone. Best score is in bold and second best is is underlined.
@@ Methods
Table 3: LiDAR segmentation results (mIoU) on the SemanticSTF [33] validation set of models trained with SemanticSTF [33]. is the reproduced result with same segmentation backbone. The symbol $\ddagger$ indicates that the validation for model selection was performed on sequence 0 of SynLiDAR, rather than SemanticSTF. Best score is in bold and second best is is underlined.
@@
SemanticKITTI to SemanticSTF. According to Tab. 2, our proposed method showed a significant improvement in SemanticSTF, with a +8.1 mIoU increase over the baseline and +5.6 mIoU over the recent competitor, PointDR. Our method demonstrated universally superior performance across all weather conditions, validating the validity of our choice in data distortion type and the effectiveness of our method. Our method substantially improved the performance of categories like other vehicle, motorcyclist, sidewalk, pole, and traffic sign, which had lower baseline performance, increasing their class IoU by about +5 to +10. Additionally, it significantly enhanced the mIoU for cars and persons by +19.0 and +9.6, respectively, compared to the baseline, which is crucial for safety in driving environments. More details are available in the supplementary material.
Also in Tab. 2, examining the performance improvement per weather condition, we note that the most significant increase in mIoU was in rain, with +7.9 mIoU. This result is indicative of our SJ augmentation effectively reflecting geometric perturbations caused by raindrops. This aspect will be further explored in the ablation study. Additionally, in snow, our method showed a remarkable increase of +7.8 mIoU, and in dense fog (+5.3 mIoU) and light fog (+7.4 mIoU), our methodology clearly outperformed the baseline. This indicates that our data augmentation effectively models the adverse impacts of weather on LiDAR semantic segmentation models. Detailed analysis of how our method enhances robustness against each weather condition, and the factors contributing to the improved performance, will be discussed in the ablation study.
@@
Figure 5: Qualitative results of our method from validation set of SemanticSTF. All models are trained in train set of SemanticKITTI.
@@
SynLiDAR to SemanticSTF. In our study, we validated the results periodically using the validation sets provided within the SemanticKITTI dataset during training. For SynLiDAR, where no specific validation set was available, sequence 0 was treated as the validation set. The best-performing models during these validation checkpoints were selected for reporting.
@@
In contrast, PointDR demonstrated an unrealistic scenario in its baseline code, using SemanticSTF for validation during training and reporting only the highest-performing results. For a fair comparison using PointDR’s criteria, our method shows a performance increase of +0.04 mIoU over PointDR. Using our more realistic protocol, our method demonstrates an even greater performance improvement of +0.32 mIoU against PointDR. Detailed results are shown in Tab. 3.
5.3 Ablation Study
@@
Analysis of Proposed Component Methods. As shown in As Tab. 4, when ASJ was added to the baseline, we observed an improvement in robustness across all adverse weather conditions (increases of +2.6 mIoU in dense fog, +5.3 mIoU in light fog, +8.1 mIoU in rain, and +6.3 mIoU in snow), albeit with a -1.8 mIoU decrease in clean data performance. The introduction of DSJ further enhanced performance in dense fog (+4.1 mIoU), light fog (+6.3 mIoU), rain (+9.3 mIoU) snow(+4.6 mIoU) and an overall performance increase of +6.2 mIoU. Our Selective Jittering method’s strong performance in rain condition showcases its efficacy in rainy weather, where it considers geometric perturbations for selected beams instead of all.
@@
The application of Range Jittering showed improvements in dense fog(+3.7 mIoU), light fog (+6.9 mIoU), rain(+6.0 mIoU), snow (+8.2 mIoU), and an overall performance increase of +7.3 mIoU, indicating the importance of simulating range distortion due to adverse weather as discussed in [10]. As seen earlier, in rainy weather, jittering only some beams aids in enhancing robustness. However, performing Range Jittering on the remaining original points leads to all points being jittered, reducing performance in rain condition. Additionally, the improvement in snow conditions by using Range Jittering suggests that perturbing as many beams as possible is more effective for robustness enhancement in snowy weather.
@@
The addition of LPD resulted in improvements in dense fog (+5.3 mIoU), light fog (+7.4 mIoU), rain (+7.9 mIoU), snow (+7.8 mIoU), and an overall increase of +8.1 mIoU against to the baseline. The substantial performance enhancements observed in conditions of dense and light fog indicate that the LPD effectively represents occlusion from foggy weather as intended. The performance reduction in snow conditions compared to the use of ASJ, DSJ, and Range Jittering (-0.4 mIoU) is likely due to LPD not exclusively targeting the ground for point dropout. LPD applies the point drop across all points, as shown in Fig. 6. This represents a limitation of LPD, necessitating the development of alternative methods to address this issue.
@@
Table 4: Experiments on component of our methods with SemanticKITTI as source and SemanticSTF as target. The values in parentheses indicate the performance improvement or decrease over the baseline model.

### v2

@@ Abstract
Existing LiDAR semantic segmentation methods often struggle with performance declines in adverse weather conditions. Previous research has addressed this issue by simulating adverse weather or employing universal data augmentation during training. However, these methods lack a detailed analysis and understanding of how adverse weather negatively affects LiDAR semantic segmentation performance. Motivated by this issue, we identified key factors of adverse weather and conducted a toy experiment to pinpoint the main causes of performance degradation: (1) Geometric perturbation due to refraction caused by fog or droplets in the air and (2) Point drop due to energy absorption and occlusions. Based on these findings, we propose new strategic data augmentation techniques. First, we introduced a Selective Jittering (SJ) that jitters points in the random range of depth (or angle) to mimic geometric perturbation. Additionally, we developed a Learnable Point Drop (LPD) to learn vulnerable erase patterns with a Deep Q-Learning Network to approximate the point drop phenomenon from adverse weather conditions. Without precise weather simulation, these techniques strengthen the LiDAR semantic segmentation model by exposing it to vulnerable conditions identified by our data-centric analysis. Experimental results confirmed the suitability of the proposed data augmentation methods for enhancing robustness against adverse weather conditions. Our method attains a remarkable 39.5 mIoU on the SemanticKITTI-to-SemanticSTF benchmark, surpassing the previous state-of-the-art by over 5.4%p, tripling the improvement over the baseline compared to previous methods achieved. Our code will be released at https://github.com/engineerJPark/LiDAR-DataAug4Weather.
@@ 1 Introduction
LiDAR semantic segmentation is a fundamental task of 3D scene understanding, particularly in safety-critical applications like autonomous driving. However, existing LiDAR segmentation models [5, 1, 14, 28, 21, 37] commonly lack robustness, showing significant performance degradation under adverse weather conditions such as snow, fog, rain, or wet surfaces.
@@
To address this issue, recent studies [11, 33, 34] have introduced corruption benchmarks for adverse weather and proposed effective techniques for robust LiDAR segmentation against corruption. These robust methods are primarily divided into task-agnostic and simulation-based approaches. The task-agnostic approach [33, 11, 25] employs general machine-learning techniques for robustness without explicitly addressing LiDAR corruption caused by adverse weather. The simulation-based approach [8, 3, 23, 7, 9, 35, 26] artificially synthesizes weather-specific data using physical equations for training. However, these efforts focus on detection tasks and consider only a single type of weather at a time.
@@
While the simulation-based approach leverages the intrinsic properties of LiDAR scans under target adverse weather conditions, simulating every weather type at all possible severities is impractical and often inaccurate. Instead, we adopt a data-centric perspective to analyze corrupted LiDAR data. For example, distortions caused by rain, snow, and fog often create similar patterns, as simulated in [9]. That is, all those adverse weather shows point drop patterns due to attenuation or occlusion by droplet [9, 7, 6]. While the simulation-based approach requires explicit modeling of the LiDAR for each weather condition, our data-centric approach potentially addresses complex distortions with a few unified patterns. Through existing research and our comprehensive analysis, we find that various adverse weather conditions create similar distortion patterns in LiDAR data. Based on our analysis, we categorize these distortions into two types: (1) geometric perturbation and (2) point drop. Our toy experiments demonstrate that these distortions are closely related to segmentation performance degradation.
@@
Rooted by the toy experiment, we introduce two novel and strategic data augmentation methods tailored to the LiDAR distortion caused by adverse weather. By incorporating these augmentations during training, we aim to bolster the model’s robustness for each distortion type. For geometric perturbation, we introduce Selective Jittering (SJ). This method applies Additive Gaussian Noise (AGN) to alter the XYZ-coordinate and intensity within a selective local region. To handle a point drop pattern, we developed a Learnable Point Drop (LPD) that employs a Deep Q-Network (DQN) to strategically remove points. Our augmentation strategies are inspired by existing studies [9, 16, 27], which shows geometric perturbation typically involves small, random alterations to the original points. This insight led to our choice of jittering as a suitable technique for addressing the geometric perturbation. Also, our augmentation strategy for point drop is motivated by existing studies [8, 9, 7] which state that intensity and depth of LiDAR beams play a significant role. This prompts us to use a DQN to detect point drops and adapt to such point drop patterns.
@@
Finally, we apply the proposed augmentations when training the LiDAR segmentation model with a clean source dataset. Experimental results demonstrate that our method achieves new state-of-the-art performance on SemanticSTF, a real adverse weather dataset. Remarkably, we achieved an impressive +8.1 mIoU gain over the baseline model, MinkowskiNet. This is a notable improvement, nearly tripling the +2.5 mIoU improvement from previous work [33]. Additionally, we observed performance gains of +7.8 and +10.3 mIoU for CENet and SPVCNN, respectively, in SemanticSTF. Our method also shows performance gains in SemanticKITTI-C, with +3.9, +0.4, and +5.6 mIoU for CENet, SPVCNN, and MinkowskiNet, respectively. Furthermore, our method outperforms recent domain generalization methods, achieving an increase of +4.9 mIoU.
@@ 2.1 LiDAR Semantic Segmentation
Voxel-based methods [5, 37, 14] perform efficient computations by dividing 3D space into a voxel grid and aggregating points within the same voxel. MinkowskiNet [5] voxelizes LiDAR points with a cubic grid and applies sparse convolution. Cylinder3D [37] proposes cylindrical partitions, reflecting LiDAR’s characteristic that the density of points depends on the distance. SphereFormer [14] utilizes a radial window and transformer architecture to aggregate long-range information and improve performance. Voxel-based methods achieve a balance between reasonable inference time and commendable segmentation performance.
@@ 2.3 LiDAR under Adverse Weather Conditions
Robustness under harsh conditions is crucial in safety-critical applications. Adverse weather substantially degrade performance in real-world outdoor autonomous driving. Thus, there are several attempts to develop weather-robust models in fields such as 2D segmentation [15, 17], 3D detection [8, 7, 11], and 3D segmentation [33, 11]. Simulation-based approaches [8, 7] artificially synthesize data for single weather conditions through physical modeling and utilize it for training. We differ from these in that we do not model specific weather conditions explicitly. Recently proposed task-agnostic approaches [33, 11, 25] consider multiple weather conditions at once. However, they use general machine-learning techniques (such as teacher-student framework and feature prototype) to achieve robustness rather than specifically tackle LiDAR corruption caused by adverse weather. We differ from these in that we propose augmentation methods specifically tailored for adverse weather conditions based on the analysis of performance degradation in LiDAR data.
@@ Methods
Based on previous research findings mentioned in Section 3, the types of distortions in LiDAR point clouds caused by adverse weather conditions converge into a set of common distortions. Therefore, from a data-centric perspective, the issues that we need to consider can be summarized as follows: (D1) Point drop, (D2) Occlusions, (D3) Geometric perturbation, and (D4) Intensity distortion. Here, we aim to identify which distortion types negatively impact performance. To achieve this, we have generated four distortion types of toy synthetic data through the SemanticKITTI validation set.
@@ 4.1 Selective Jittering
Depth-Selective Jittering (DSJ). DSJ employs a single scan to add Gaussian noise to the XYZ coordinates and the intensity value of points below a randomly selected depth range.
@@ 4.2 Learnable Point Drop
In Comparison with random drop, random drop uniformly reduces points across all depths. Consequently, this approach is not suitable for simulating point drops caused by fog, as it lacks the depth-specific characteristics of fog-related distortions. Furthermore, randomly generated point drop scans may not achieve the severity of adverse weather conditions necessary to drop critical points that could deceive the model. A detailed comparison will be conducted in the supplementary material.
@@ 5.1 Experimental Setup
Dataset and Evaluation Metrics. To assess robustness, we trained our model on 19 classes from SemanticKITTI and then validated its performance on validation set of SemanticSTF as same as [33]. The experiment process involves training on clean data followed by evaluating robustness in adverse conditions, akin to existing research [11, 33, 34]. SemanticSTF is a dataset gathered under real-world adverse conditions such as rain, fog, and snow. Following the approach in [33], invalid data in SemanticSTF are mapped to an “ignore label”. We utilized MinkowskiNet [5] for validation purposes due to its well-known robustness [11, 34]. Our evaluation metrics include the Intersection over Union (IoU) for each class and the mean IoU (mIoU) across all classes.
@@ 5.2 Main Results
Table 2: LiDAR segmentation results (mIoU) on the SemanticSTF validation set of models trained with (a) SemanticKITTI and (b) SynLiDAR [32]. D-fog and L-fog denote dense fog and light fog weather conditions in all experiments. is the reproduced result with the same segmentation backbone. The symbol $\ddagger$ indicates that the validation for model selection was performed on sequence 0 of SynLiDAR, rather than SemanticSTF. Best score is in bold and second best is underlined.
[…]
@@ Methods
[…]
@@
SemanticKITTI to SemanticSTF. According to Tab. 2 (a), our proposed method showed a significant improvement in SemanticSTF, with a +8.1 mIoU increase over the baseline and +5.6 mIoU over the recent competitor, PointDR. Our method demonstrated universally superior performance across all weather conditions, validating the validity of our choice in data distortion type and the effectiveness of our method. Our method substantially improved the performance of categories like other vehicle, motorcyclist, sidewalk, pole, and traffic sign, which had lower baseline performance, increasing their class IoU by about +5 to +10. Additionally, it significantly enhanced the mIoU for cars and persons by +19.0 and +9.6, respectively, compared to the baseline, which is crucial for safety in driving environments. More details are available in the supplementary material.
Also, in Tab. 2 (a), examining the performance improvement per weather condition, we note that the most significant increase in mIoU was in rain, with +7.9 mIoU. This result is indicative of our SJ augmentation effectively reflecting geometric perturbations caused by raindrops. This aspect will be further explored in the ablation study. Additionally, in snow, our method showed a remarkable increase of +7.8 mIoU, and in dense fog (+5.3 mIoU) and light fog (+7.4 mIoU), our methodology clearly outperformed the baseline. This indicates that our data augmentation effectively models the adverse impacts of weather on LiDAR semantic segmentation models. Detailed analysis of how our method enhances robustness against each weather condition, and the factors contributing to the improved performance, will be discussed in the ablation study.
@@
Figure 4: Qualitative results of our method from validation set of SemanticSTF. All models are trained in train set of SemanticKITTI. Green (or red) points indicate correct (or incorrect) predictions, showing our predictions are more accurate than PointDR in all weather, especially on road and sidewalk.
@@
SynLiDAR to SemanticSTF. For SynLiDAR, unlike SemanticKITTI, where no specific validation set was available, sequence 0 was treated as the validation set. The best-performing models during these validation checkpoints were selected for reporting.
@@
In contrast, PointDR demonstrated an unrealistic scenario in its baseline code, using SemanticSTF for validation during training and reporting only the highest-performing results. For a fair comparison using PointDR’s criteria, our method shows a performance increase of +0.04 mIoU over PointDR. Using our more realistic protocol, our method demonstrates an even greater performance improvement of +0.32 mIoU against PointDR. Detailed results are shown in Tab. 2 (b).
5.3 Additional Experiments
@@
Analysis of Proposed Component Methods. As shown in As Tab. 3, when ASJ was added to the baseline, we observed an improvement in robustness across all adverse weather conditions (increases of +2.6 mIoU in dense fog, +5.3 mIoU in light fog, +8.1 mIoU in rain, and +6.3 mIoU in snow), albeit with a -1.8 mIoU decrease in clean data performance. The introduction of DSJ further enhanced performance in dense fog (+4.1 mIoU), light fog (+6.3 mIoU), rain (+9.3 mIoU), snow (+4.6 mIoU), and an overall performance increase of +6.2 mIoU. Our Selective Jittering method’s strong performance in rain conditions showcases its efficacy in rainy weather, where it considers geometric perturbations for selected beams instead of all.
@@
The application of Range Jittering showed improvements in dense fog(+3.7 mIoU), light fog (+6.9 mIoU), rain(+6.0 mIoU), snow (+8.2 mIoU), and an overall performance increase of +7.3 mIoU, indicating the importance of simulating range distortion due to adverse weather as discussed in [9]. As seen earlier, in rainy weather, jittering only some beams aids in enhancing robustness. However, performing Range Jittering on the remaining original points leads to all points being jittered, reducing performance in rain conditions. Additionally, the improvement in snow conditions by using Range Jittering suggests that perturbing as many beams as possible is more effective for robustness enhancement in snowy weather.
@@
The addition of LPD resulted in improvements in dense fog (+5.3 mIoU), light fog (+7.4 mIoU), rain (+7.9 mIoU), snow (+7.8 mIoU), and an overall increase of +8.1 mIoU against to the baseline. The substantial performance enhancements observed in conditions of dense and light fog indicate that the LPD effectively represents occlusion from foggy weather as intended. The performance reduction in snow conditions compared to the use of ASJ, DSJ, and Range Jittering (-0.4 mIoU) is likely due to LPD not exclusively targeting the frozen or wet ground for point dropout. LPD applies the point drop across all points, as shown in Fig. 5. This represents a limitation of LPD, necessitating the development of alternative methods to address this issue.
@@
Table 3: Experiments on the components of our methods with SemanticKITTI as source and SemanticSTF as target. The values in parentheses indicate the performance improvement or decrease over the baseline model.